OpenStudy (anonymous):

Turn 2x+3y=-9 into slope-intercept form.

OpenStudy (anonymous):

@igreen

OpenStudy (anonymous):

2x + 3y - 9 We have to isolate y on one side, so subtract 2x from both sides. 2x + 3y - 9 -2x -2x 3y = -2x - 9 Divide each side by three to get y by itself. 3y/3 = -2x/3 - 9/3 y = -2/3x - 3 <----- This is the answer. (:

OpenStudy (anonymous):

THank you so much!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

You're welcome! (:

OpenStudy (anonymous):

Can you help me with one more?